# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ 6 Continuous Nets Sample Complexity
So far, we focused on quantized uniform priors. However, per-layer continuous spherical priors (e.g., Gaussian) are also quite common in practice and theory. Therefore, in this section, we show how to extend our results beyond the quantized case into a continuous setting, for the special case of two-layer NNs without bias and with the leaky rectifier linear unit (Maas et al., 2013) activation function. Formally, let $h_{\bm{\theta}},h_{\bm{\theta}^{\star}}$ be fully connected (Definition 5.1) two layer NNs with input dimension $d_{0}$, output dimension $d_{2}=1$ and hidden layer dimensions $d_{1}$ and $d_{1}^{\star}$, respectively.
$h_{\bm{\theta}}\left(\mathbf{x}\right)=\operatorname{sign}\left(\mathbf{W}^{\left(2\right)}\sigma\left(\mathbf{W}^{\left(1\right)}\mathbf{x}\right)\right)$
$h_{\bm{\theta}^{\star}}\left(\mathbf{x}\right)=\operatorname{sign}\left(\mathbf{W}^{\left(2\right)}_{\star}\sigma\left(\mathbf{W}^{\left(1\right)}_{\star}\mathbf{x}\right)\right)$
@@
and $\sigma\left(\cdot\right)$ is the common leaky rectifier linear unit (LReLU, Maas et al. (2013)) with parameter $\rho\notin\{0,1\}$
As in the previous sections, our goal is to obtain generalization guarantees by lower bounding $\tilde{p}$ and then combining this results with Theorem 3.2. To this end, we first need to define some prior on G&C hypothesis sampling.
@@
Suppose that the weights of $h_{\bm{\theta}}$ are random such that each row of the first layer, $\mathbf{w}_{i}$, is independently sampled from a uniform distribution on the unit sphere $\mathbb{S}^{d_{0}-1}$, and the second layer $\mathbf{W}^{\left(2\right)}$ is sampled uniformly [^8] Sampling $\mathbf{W}^{\left(2\right)}$ from the unit sphere is equivalent to sampling it from a Gaussian distribution. In fact, any spherically symmetric distribution in $\mathbb{R}^{d_{1}}$ will suffice, as it amounts to scaling of the output without affecting the classification. from $\mathbb{S}^{d_{1}-1}$. Both $\mathbf{w}_{i}$ and $\mathbf{W}^{\left(2\right)}$ are independent of the teacher and data.
@@
To extend this notion of teacher-equivalence from the quantized setting to the continuous setting, we assume an “angular margin" exists between the training set $\mathcal{S}$ and the teacher, similarly to Soudry and Hoffer (2017).
@@
Let $\alpha<\beta\in\left(0,\frac{\pi}{2}\right)$. There exist $\lambda\in\left(0,1\right)$ such that, with probability at least $1-\lambda$ over $\mathcal{S}\sim\mathcal{D}^{N}$, $\mathcal{S}$ has first layer angular margin $\alpha$ (Definition 6.2) and second layer angular margin $\beta$ (Definition 6.3).
@@
Note that $\alpha$ is the minimal margin of all hidden neurons, while $\beta$ is the margin of the single network output. Thus, intuitively, $\beta$ is usually larger than $\alpha$. For Gaussian data, we show empirically in Figure 3 in Appendix C, that the assumption $\beta>\alpha$ holds with high probability.
This assumption allows us to extend the results from the previous section to a continuous setting.
Definition 6.6.
For some random hypothesis $h$ from the prior $\mathcal{P}$, the interpolation probability is defined as
$\hat{p}\triangleq\mathbb{P}_{h\sim\mathcal{P}}\left(\mathcal{L}_{\mathcal{S}}(h)=0\right)$
@@
Assume that $\hat{p}<\frac{1}{2}$ and $d_{0}\gg d_{1}^{\star}\gg 1$  [^9] This assumption is used to simplify the bound. A non-asymptotic version that does not require this assumption is in Appendix C.. Then under Assumption 6.4, for any $\epsilon,\delta\in\left(0,1\right)$ we have
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)\geq 1-\delta\ -\lambda,$
@@
$\frac{\hat{C}^{\mathrm{cont}}+2\log\left(\hat{C}^{\mathrm{cont}}\right)+4\log\left(\frac{8}{\delta}\right)}{\epsilon}\,,$
@@
The full proof is in Appendix C.
@@
Since Assumption 6.4 gives margin only in high probability, we use a different generalization guarantee than the one in Theorem 3.2 which takes into account the interpolation probability $\hat{p}$, instead of $\tilde{p}$. See Appendix A.6 for more details.
@@ 7 Related work
Random interpolating neural networks. Chiang et al. (2023) empirically studied a few gradient-free algorithms for the optimization of neural networks, suggesting that generalization is not necessarily due to the properties of the SGD. For the G&C algorithm, they also empirically investigated the effect of the loss on the generalization. We focus on the 0-1, rather than some surrogate loss function. Chiang et al. (2023) focused only on small-scale datasets since it is not possible to run their experiments when the training set is larger than a few dozen samples. Theisen et al. (2021) provided a theoretical analysis of random interpolating linear models. In contrast, our work focuses on deeper models with non-linear activation functions, and our teacher assumption relies on depth, such that it is possible to have many zeroed-out hidden neurons and obtain generalization guarantees that stem from the redundancy in parameterized deep networks. Valle-Perez et al. (2019); Mingard et al. (2021) used a connection between neural networks and Gaussian processes to model random interpolating networks, which usually requires the width of the network to be infinite, in contrast with our finite width setting.
Redundancy in neural networks. The Lottery Ticket Hypothesis (Frankle and Carbin, 2019) suggests that for any random neural network, there exists a sparse sub-network capable of competitive generalization with the original. A related hypothesis was conjectured by Ramanujan et al. (2020) and later proven by Malach et al. (2020) and states that it is possible to prune an initialized neural network without significantly affecting the obtained predictions. In contrast, our work, while also focusing on random neural networks, is oriented towards providing assurances regarding the generalization of these random networks, when conditioned on the event of perfectly classifying the training data. Also, we operate under the assumption of a narrow teacher, which is sparse in the number of neurons — in contrast to the sub-network in the Lottery Ticket Hypothesis which is sparse in the number of weights.
@@ 8 Discussion
Implications of narrowness. Assuming a narrow teacher may sometimes limit the class of possible target functions. For example, in a fully connected NN, if the first hidden layer width is smaller than the input width $\left(d_{0}>d_{1}^{\star}\right)$ then the teacher NN must be constant on the $d_{0}-d_{1}^{*}$ nullspace of the first weight layer. However, this is not an issue for a convolutional NN, since typically the number of input channels is small (e.g., $c_{0}=3$ for RGB images), and so $c_{0}<c_{1}^{*}$ is reasonable. In any case, to represent potentially complex target functions with narrow NNs we require depth (Kidger and Lyons, 2020) — both for student and teacher. This might suggest another reason why deep architectures are more useful in realistic settings.
@@
Quantized models. Our results in section 5 rely on NNs to have a quantized numerical format. The sample complexity we obtain is a product of the parameter count of the model (mainly the teacher) and the quantization bits. The simplicity of the bound allows it to be applied to other architectures as well. For example, we can add pooling layers to the CNN models we analyzed. Also, using the same considerations, we may obtain similar results for multi-head attention layers. Moreover, using quantization-aware methods (Hubara et al., 2018) one can reduce numerical precision to improve the bound. For example, using 2bits weights (and activations) in ResNet50 on ImageNet results in only 0.5% degradation in accuracy (Liu et al., 2022)). Such quantization approaches are common in compression-type bounds (e.g. Lotfi et al. (2022)). However, there the goal is to compress the learned model (the student), while here we only need some compressed model to exist (i.e. the narrow teacher).
@@
Beyond interpolators. In this work we focus for simplicity on NN interpolators, but in many scenarios, the training loss does not reach zero. It should be straightforward to generalize our results to this non-realizable case, where the teacher NN does not reach zero loss (i.e., there is some irreducible error). Additionally, the implicit bias of SGD is partially known in some cases (e.g., Lyu and Li (2020)), so it would be interesting to understand the generalization of a typical NN interpolator that also obeys this implicit bias.
Acknowledgements
@@
Appendix A Generalization Results

### v2

@@ 5 Continuous Nets Sample Complexity
So far, we focused on quantized uniform priors. However, per-layer continuous spherical priors (e.g., Gaussian) are also quite common in practice and theory. Therefore, in this section, we show how to extend our results beyond the quantized case into a continuous setting, for the special case of two-layer NNs without bias and with the leaky rectifier linear unit (LReLU, Maas et al. (2013)) activation function. Formally, let $h_{\boldsymbol{\theta}},h_{\boldsymbol{\theta}^{\star}}$ be fully connected (Definition 4.1) two layer NNs with input dimension $d_{0}$, output dimension $d_{2}=1$ and hidden layer dimensions $d_{1}$ and $d_{1}^{\star}$, respectively. Explicitly:
$h_{\boldsymbol{\theta}}\left(\mathbf{x}\right)=\operatorname{sign}\left(\mathbf{W}^{\left(2\right)}\sigma\left(\mathbf{W}^{\left(1\right)}\mathbf{x}\right)\right)$
$h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{x}\right)=\operatorname{sign}\left(\mathbf{W}^{\left(2\right)}_{\star}\sigma\left(\mathbf{W}^{\left(1\right)}_{\star}\mathbf{x}\right)\right)$
@@
and $\sigma\left(\cdot\right)$ is the common LReLU with parameter $\rho\notin\{0,1\}$.
As in the previous sections, our goal is to obtain generalization guarantees by lower bounding $\tilde{p}$ and then combining this results with Lemma 3.2. To this end, we first need to define some prior on the hypotheses.
@@
Suppose that the weights of $h_{\boldsymbol{\theta}}$ are random such that each row of the first layer, $\mathbf{w}_{i}$, is independently sampled from a uniform distribution on the unit sphere $\mathbb{S}^{d_{0}-1}$, and the second layer $\mathbf{W}^{\left(2\right)}$ is sampled uniformly [^4] Sampling $\mathbf{W}^{\left(2\right)}$ from the unit sphere is equivalent to sampling it from a Gaussian distribution. In fact, any spherically symmetric distribution in $\mathbb{R}^{d_{1}}$ will suffice, as it amounts to scaling of the output without affecting the classification. from $\mathbb{S}^{d_{1}-1}$. Both $\mathbf{W}^{\left((1)\right)}$ and $\mathbf{W}^{\left(2\right)}$ are independent of the teacher and data.
@@
To extend the notion of teacher-equivalence from the quantized setting to the continuous setting, we assume an “angular margin” exists between the training set $\mathcal{S}$ and the teacher, similarly to Soudry & Hoffer (2017).
@@
Let $\alpha<\beta\in\left(0,\frac{\pi}{2}\right)$. There exists $\lambda\in\left(0,1\right)$ such that with probability at least $1-\lambda$ over $\mathcal{S}\sim\mathcal{D}^{N}$, $\mathcal{S}$ has first layer angular margin $\alpha$ (Definition 5.2) and second layer angular margin $\beta$ (Definition 5.3).
@@
Note that $\alpha$ is the minimal margin of all hidden neurons, while $\beta$ is the margin of the single network output. Thus, intuitively, $\beta$ is usually larger than $\alpha$. For Gaussian data, we show empirically in Figure 3 in Appendix D, that the assumption $\beta>\alpha$ holds with high probability.
This assumption allows us to extend the results from the previous section to a continuous setting (proof in Appendix D).
@@
Assume that $\hat{p}_{\mathcal{S}}<\frac{1}{2}$ a.s. and $d_{0}\gg d_{1}^{\star}\gg 1$  [^5] This assumption is used to simplify the bound. A non-asymptotic version that does not require this assumption is in Appendix D.. Then under Assumption 5.4, for any $\varepsilon,\delta\in\left(0,1\right)$ we have
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)\geq 1-\delta\ -\lambda,$
@@
$\frac{\hat{C}^{\mathrm{cont}}+2\log\left(\hat{C}^{\mathrm{cont}}\right)+4\log\left(\frac{8}{\delta}\right)}{\varepsilon}\,,$
@@
$\hat{C}^{\mathrm{cont}}$ $=-d_{1}^{\star}d_{0}\log\left(\sin\left(\alpha\right)\right)-d_{1}\log\left(\sin\left(\gamma\right)\right)$
$\quad+\frac{1}{2}d_{1}^{\star}\log\left(d_{0}\right)+O\left(d_{1}^{\star}+\log{\left(d_{1}\right)}\right)\,.$
@@
Since Assumption 5.4 gives a positive margin only in high probability, we use a different generalization guarantee than the one in Lemma 3.2 which takes into account the interpolation probability $\hat{p}_{\mathcal{S}}$, instead of $\tilde{p}$. See Appendix B.7 for more details.
Remark 5.8.
Although Assumption 5.4 may be natural in some cases [^6] Realistic data many times has some intrinsic margin. For example, there is a low probability for (semantic) ‘mixing’ of dogs and birds from a distribution of natural images of dogs and birds (with a single animal in every image). , in other cases the margins $\alpha,\beta$ may decay with $N$ or with the network’s dimensions. Therefore, even if the assumption holds, for the generalization bound in Theorem 5.6 to remain meaningful, we need the margin decay rate to be sufficiently slow. [^7] For example, for Gaussian data with a random teacher it is possible to derive a lower bound in high probability on the first layer angular margin $\alpha$ and show that it decays as $1/(Nd_{0}d_{1}^{\star})$, so it adds only a multiplicative $\log(Nd_{0}d_{1}^{\star})$ factor to the bound.
@@ 6 Related work
Random interpolating neural networks. Chiang et al. (2023) empirically studied a few gradient-free algorithms for the optimization of neural networks, suggesting that generalization is not necessarily due to the properties of the SGD. For the G&C algorithm, they also empirically investigated the effect of the loss on the generalization. We focus on the 0-1, rather than some surrogate loss function. Chiang et al. (2023) focused only on small-scale datasets since it is not possible to run their experiments when the training set is larger than a few dozen samples. Theisen et al. (2021) provided a theoretical analysis of random interpolating linear models. In contrast, our work focuses on deeper models with non-linear activation functions, and our teacher assumption relies on depth, such that it is possible to have many zeroed-out hidden neurons and obtain generalization guarantees that stem from the redundancy in parameterized deep networks. Valle-Perez et al. (2019); Mingard et al. (2021) used a connection between neural networks and Gaussian processes to model random interpolating networks, which usually requires the width of the network to be infinite, in contrast with our finite width setting. Teney et al. (2024) empirically investigated random fully connected neural networks without conditioning on the interpolation of a training set. They examined their spectral properties and found a bias towards simple functions (according to various metrics). Berchenko (2024) showed generalization results for uniformly sampled trees learning Boolean functions, and discuss the simplicity bias of random deep NNs. In contrast, we proved an explicit generalization bound for typical interpolating deep neural networks.
Redundancy in neural networks. The Lottery Ticket Hypothesis (Frankle & Carbin, 2019) suggests that for any random neural network, there exists a sparse sub-network capable of competitive generalization with the original. A related hypothesis was conjectured by Ramanujan et al. (2020) and later proven by Malach et al. (2020) and states that it is possible to prune an initialized neural network without significantly affecting the obtained predictions. In contrast, our work, while also focusing on random neural networks, is oriented towards providing assurances regarding the generalization of these random networks, when conditioned on the event of perfectly classifying the training data. Also, we operate under the assumption of a narrow teacher, which is sparse in the number of neurons — in contrast to the sub-network in the Lottery Ticket Hypothesis which is sparse in the number of weights.
@@ 7 Discussion
Implications of narrowness. Assuming a narrow teacher may sometimes limit the class of possible target functions. For example, in a fully connected NN, if the first hidden layer width is smaller than the input width $\left(d_{0}>d_{1}^{\star}\right)$ then the teacher NN must be constant on the $d_{0}-d_{1}^{*}$ nullspace of the first weight layer. However, this is not an issue for a convolutional NN, since typically the number of input channels is small (e.g., $c_{0}=3$ for RGB images), and so $c_{0}<c_{1}^{*}$ is reasonable. In any case, (1) many realistic datasets are low rank (Udell & Townsend, 2019; Zeno et al., 2024), and (2) in order to represent potentially complex target functions with narrow NNs we require depth (Kidger & Lyons, 2020) — both for student and teacher. This might suggest another reason why deep architectures are more useful in realistic settings.
@@
Quantized models. Our results in Section 4 rely on NNs to have a quantized weights. The sample complexity we obtain is a product of the parameter count of the model (mainly the teacher) and the quantization bits. The simplicity of the bound allows it to be applied to other architectures as well. For example, we can add pooling layers to the CNN models we analyzed. Also, using the same considerations, we may obtain similar results for multi-head attention layers when the student has redundant heads and no LayerNorm. Moreover, using quantization-aware methods (Hubara et al., 2018) one can reduce numerical precision to improve the bound. For example, using 2bits weights (and activations) in ResNet50 on ImageNet results in only 0.5% degradation in accuracy (Liu et al., 2022a)). Such quantization approaches are common in compression-type bounds (e.g. Lotfi et al. (2022)). However, there the goal is to compress the learned model (the student), while here we only need some compressed model to exist (i.e. the narrow teacher).
@@
Beyond interpolators. In this work we focus for simplicity on NN interpolators, but in many scenarios, the training loss does not reach zero. In Appendix B.4 we show how to generalize our results to this non-realizable case, where the teacher NN does not reach zero training error (i.e., there is some irreducible error). Unfortunately, the use of the non-realizable generalization bound for finite hypothesis classes introduce a quadratic dependence of the sample complexity on the generalization error bound $\varepsilon$. It is interesting to see if this quadratic dependence can be improved.
Does posterior sampling bias towards sparse representations? Our results indicate that posterior sampling biases NNs towards sparse representations. To examine whether this also happens in other models, we empirically examined posterior sampling in a sparse regression setting for linear diagonal networks (a common theoretical model, e.g. Woodworth et al. (2020); Moroshko et al. (2020)) with a Gaussian weights prior, where the ground truth has a single nonzero component. In that case, we do get a bias toward a sparser predictor, but only with sufficient depth. Specifically, we found that a small depth (2 or 3) does not help much to improve generalization compared to depth 1. However, larger depth did seem to help significantly. In contrast, the NN’s result in this paper is different, since there depth is not necessary to obtain good generalization results (as our results hold even with depth 2).
Parameterization and minimum description length. The results in this paper rely heavily on the choice of parameterization of the hypotheses. Specifically, as discussed in Section 3, we can obtain results similar to Lemma 3.2 using a Minimum Description Length (MDL)/Occam’s Razor learning rule. Choosing a different mapping from parameters (or descriptions) to hypotheses may not benefit from the redundancy and will result in different generalization bounds which may be worse and even trivial. An example for such parameterization and further discussion are presented in Appendix A.
Relation to SGD. As we have mentioned in Remark 4.4, the volume of ‘bad’ interpolators is exponentially decaying in the size of the training set. Therefore, algorithms with bad generalization properties must have significant probability to sample this ‘small’ subset of hypotheses. Therefore, it would be interesting to know when practical training algorithms are biased towards this ‘bad’ domain. For example, the implicit bias of SGD is partially known in some cases (e.g., Lyu & Li (2020)), so it would be interesting to understand the generalization of typical NN interpolators that also obey this implicit bias.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
@@
Appendix A Redundancy and Description Length
In this section, we want to give a broader perspective on the effect of the parameterization on $\tilde{p}=\mathbb{P}_{h\sim\mathcal{P}}\left(h\equiv h^{\star}\right)$. As we saw, the parameterization controls the generalization behavior of G&C and posterior sampling, and how a seemingly uniform prior over parameters can induce a rich prior over hypothesis via redundancy. We first consider a conceptually very direct parameterization.
How parameter redundancy works: a minimal example. Consider a predictor $h_{\sigma}$ specified by a description language (i.e. parameterization) using bit strings $\sigma$, where descriptions end with the string “END” (encoded in bits), and this string never appears elsewhere in the description [^8] This can be a Turing complete programming language, or perhaps better to think of simpler descriptions such as boolean formulas or strings encoding decision trees.. We will consider students that use fixed-length descriptions, i.e., $\sigma\in\{0,1\}^{C}$, where we only consider the description up to the first “END”, i.e. $h_{\sigma}=h_{\tilde{\sigma}}$ where $\tilde{\sigma}$ is a prefix of $\sigma$ ending with the first “END”. [^9] If $\sigma$ does not contain “END”, or $\tilde{\sigma}$ is not a valid description, we can set $h_{\sigma}$ to the constant 0 predictor. Let us now consider a uniform prior over this parameterization, i.e. $\mathcal{P}(\sigma)=2^{-C}$ for all $\sigma\in\{0,1\}^{C}$. Although this prior is uniform over parameterizations, it is easy to see that it induces a highly non-uniform prior over predictors: for a predictor $h_{\tau}$ with a short description $\tau\in\{0,1\}^{{\left\lvert{\tau}\right\rvert}}$, ${\left\lvert{\tau}\right\rvert}\ll C$, it’s $C$-bit description is highly redundant, as the final $C-{\left\lvert{\tau}\right\rvert}$ bits can be set arbitrarily, as long as the first bits are $\tau$, and so $\mathcal{P}(h_{\tau})\geq 2^{-{\left\lvert{\tau}\right\rvert}}$. For a teacher with a short description $h^{\star}=h_{\tau^{\star}}$, $\tau^{\star}\in\{0,1\}^{C^{\star}}$, we thus have $-\log\tilde{p}\leq C^{\star}$. The sample complexity of learning a small teacher thus depends only on the complexity of the teacher and not the complexity of the (possibly much larger) student.
Why would MDL fail here? Returning to the MDL / Occam principle mentioned at the end of Section 3, it is important to note that applying this principle to the distribution over parameters would not work here. Consider the MDL / Occam / MAP rule which selects the most likely interpolating parameters $\arg\max_{\mathcal{L}_{\mathcal{S}}(h_{\sigma})}\mathcal{P}(\sigma)$. This rule would be useless here, since for all $\sigma$ in our parameter space $\{0,1\}^{C}$ have the same prior probability, and its associated sample complexity would be $O(-\log\mathcal{P}(\sigma^{\star}))=C$ (not $\mathcal{P}(h_{\sigma^{\star}})$ !), where $\sigma^{\star}$ is a description of $h^{\star}$ in our parameter space (e.g. $\sigma^{\star}=\tau^{\star}+$‘0’$\cdot(C-C^{\star})$). The important distinction is that Occam considers the probability mass function over parameters (roughly, the density under a specific parameterization or base measure). On the other hand, even if we sample $h_{\sigma}$ by sampling parameters $\sigma$, posterior sampling (i.e. G&C) can be thought of directly in terms of the distribution over hypotheses.
When would redundancy fail? To see how changing the parameterization can change the induced distribution over hypothesis, consider instead a non-redundant parameterization, where we take $\sigma$ to be uniform over valid descriptions, of length at most $C$, ending with “END” (i.e. strings ending with “END” and that do not otherwise contain “END”). The induced prior over hypothesis is now uniform [^10] We might still have further redundancies in that multiple valid descriptions can describe the same function, again introducing non-uniformity. Consider here a non-redundant description language, e.g. a read-once branching program. and even using posterior sampling/G&C would have sample complexity determined by the complexity of the student, rather than the teacher.
When would MDL succeed? Finally, we note that we could of course learn a short teacher with an Occam rule that is explicitly biased towards short descriptions, e.g. using the Kraft prior [^11] We again absorb any remaining probability in the constant zero predictor. Here we are thinking of an unbounded student and the parameter space being any string ending with the first occurrence of “END”. We can also of course bound the size of the student. $\mathcal{P}(\sigma)=2^{-{\left\lvert{\sigma}\right\rvert}}$. Using such a prior, the Occam rule also enjoys a sample complexity of $O(C^{\star})$ that depends only on the length of the teacher. But here the prior over parameters is non-uniform and explicitly biased, and our interest in this paper is in how seemingly uniform priors over parameters can induce non-uniform priors and generalization over predictors due to the choice of parameterization.
Appendix B Generalization Results